OpenStudy (anonymous):

In which of the following groups of fractions can all the fractions be renamed as a whole number of hundredths?

OpenStudy (anonymous):

\[a. 3/2 , 3/4, 3/5 \] \[b. 7/10,7/11,7/12\] \[c. 2/5,2/6,2/8\] \[d. 11/5,11/10,11/20\]

OpenStudy (unklerhaukus):

\[a.\]\[\frac32=\frac{3\times50}{2\times50}=\frac{150}{100}\]\[\frac34=\frac{3\times25}{4\times25}=\frac{75}{100}\]\[\frac35=\frac{3\times20}{5\times20}=\frac{60}{100}\]